In right triangle ABC, with right angle ACB, AC=8 inches and BC= 6 inches. Point X is on segment AC, equidistant from A and B. Find CX. I need some help with this. Thanks!!!
Answer by scott8148(5880)

let x=length of CX

triangle AXB is isosceles

by Pythagoras, 6^2 + x^2 = (8-x)^2

36 + x^2 = 64 -16x + x^2

subtracting x^2 __ 36 = 64 - 16x

adding 16x-36 __ 16x = 28

dividing by 16 __ x = 1.75